(a) Calculate the capacitance of a parallel-plate capacitor whose plates are 20 cm x 30 cm and are separated by a 1.0-mm air gap.


(b) what is the charge on each plate if a 12-V battery is connected across the two plates?


(c) What is the electric field between the plates?


1
Expert's answer
2022-03-27T10:56:04-0400

(a)

"C=\frac{\epsilon_0 A}{d}\\
=\frac{8.85*10^{-12}*0.2*0.3}{0.001}=5.31*10^{-10}\:\rm F"

(b)

"q=CV\\
=5.31*10^{-10}*12=6.37*10^{-9}\:\rm C"

(c)

"E=\frac{V}{d}\\
=\frac{12}{0.001}=12000\:\rm V/m"